# Low-Rank Online Dynamic Assortment with Dual Contextual Information

### v1

@@ 2 Problem Formulation
In this section, we introduce the mathematical formulation of the LOAD problem. At time $t$, a user with feature vector $q_{t}\in\operatorname*{\mathbb{R}}^{d_{1}}$ arrives at the platform. Assume there are $N$ products, each with item feature vector $p_{i}\in\operatorname*{\mathbb{R}}^{d_{2}}$ for $i\in[N]=\{1,\ldots,N\}$. The platform offers an assortment $S_{t}$ of size at most $K$ from the catalog $[N]$. In other words, $S_{t}\subset[N]$ and $|S_{t}|\leq K$. When the user is provided with an assortment, the user either chooses one of the items in the assortment or chooses not to purchase at all. To quantify this user choice, let $y_{it}\in\{0,1\}$ represent whether user $t$ chooses item $i$, for $i\in[N]$ and $t\geq 1$. In addition, denote $y_{0t}=1$ as the indicator that user $t$ does not choose any item at all. Let $i_{t}\in S_{t}$ be the item that user $t$ chooses (00 if no purchase). Note that for at each user $t$, $y_{i_{t},t}=1$ and the remaining is $y_{jt}=0$ for $j\in S_{t}\cup\{0\}\backslash\{i_{t}\}$, i.e., $\sum_{i\in S_{t}\cup\{0\}}y_{it}=1$. Denote $\textbf{y}_{t}=(y_{1t},\ldots,y_{Nt})^{\top}\in\{0,1\}^{N}$ as the choice vector of user $t$.
In assortment selection, a widely adopted choice model is the multinomial logit (MNL) model (McFadden, , 1974; Rusmevichientong et al., , 2010; Agrawal et al., , 2019). Suppose the utility of item $i$ to user $t$ is given as a random variable $X_{it}=v_{it}+\epsilon_{it}$, where $v_{it}$ is the expected utility and $\epsilon_{it}$ is a mean-zero error term.Assume $v_{0t}=0$ for identifiability. The MNL model assumes that the probability of user $t$ choosing item $i$ is given as:
@@
where $\Phi^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ is a rank-$r$ matrix, illustrated in Figure 4, with singular values $\sigma_{1}>\cdots\sigma_{r}>0$.
After the user’s purchasing decision under the MNL model, the platform earns revenue from the selected item (or zero if no purchase is made). Let $r_{i}$ denote the revenue that the platform achieves when item $i$ is sold. Given an assortment $S$, the revenue that the platform achieves from user $t$ is $\sum_{i\in S}r_{i}y_{it}$, and the expected revenue of the platform for user $t$ is:
@@
The platform aims to choose the personalized assortment for each user $t$ at time $t$ to maximize the cumulative expected revenue gained from all users. Note that the possible assortment set with constraint $K$, i.e., $\operatorname*{\mathcal{S}}=\{S\subset[N]:|S|\leq K\}$ is finite. Therefore, when $\Phi^{*}$ is known in hindsight, there exists an optimal assortment $S_{t}^{*}$ for user $t$ that maximizes the expected revenue since we can calculate the expected utility for every $S\in\operatorname*{\mathcal{S}}$. Denote the optimal assortment at time $t$ as:
@@
Note that maximizing the cumulative revenue is equivalent to minimizing the cumulative regret. At time $t$, given the current user context $q_{t}$ and historical data, including past assortments $S_{1},S_{2},\ldots,S_{t-1}$ presented to users $11,2,\ldots,t-$ 1 with respective features $q_{1},q_{2},\ldots,q_{t-1}$, and the corresponding feedback $\textbf{y}_{1},\textbf{y}_{2},\ldots,\textbf{y}_{t-1}$, the platform needs to determine the assortment $S_{t}$ for the user $t$ so that the cumulative regret could be small.
@@ 3.1 Estimation of Subspace
The first stage of the algorithm aims to acquire estimations of low-rank subspaces. Let $\Phi^{*}=U^{*}D^{*}(V^{*})^{\top}$ be the singular value decomposition of $\Phi^{*}$, where $U^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{2}},V^{*}\in\operatorname*{\mathbb{R}}^{d_{1}\times d_{1}}$ are square orthonormal matrices and $D^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ is a diagonal matrix with singular values as its diagonal components. Let $u^{*}_{(j)}$ be the $j$-th column of $U^{*}$ and $v^{*}_{(j)}$ be the $j$-th column of $V^{*}$. Consider dual contexts $p\in\operatorname*{\mathbb{R}}^{d_{2}}$ and $q\in\operatorname*{\mathbb{R}}^{d_{1}}$. The bilinear form of $p$ and $q$ with respect to $\Phi^{*}$ can be rewritten as:
@@
where $\operatorname*{\mathcal{L}}_{n}(\Phi)$ is the negative log-likelihood of the MNL choice model with $n$ samples:
@@
To solve this rank-constrained optimization, we use the Burer-Monteiro formulation (Burer and Monteiro, , 2003) to optimize over $U\in\operatorname*{\mathbb{R}}^{d_{2}\times r}$ and $V\in\operatorname*{\mathbb{R}}^{d_{1}\times r}$, where $\Phi=UV^{\top}$. We add a regularization term to enforce identifiability (Zheng and Lafferty, , 2016). This leads to our final optimization problem:
@@
$U\leftarrow U-\eta\cdot\nabla_{U}\tilde{\operatorname*{\mathcal{L}}}_{n}(U,V),\quad V\leftarrow V-\eta\cdot\nabla_{V}\tilde{\operatorname*{\mathcal{L}}}_{n}(U,V),$
where $\eta$ is the learning rate, $\nabla_{U},\nabla_{V}$ are gradients with respect to each component when other parameters are fixed. Applying these gradient descent steps alternatively, the estimated parameters are known to converge to the true parameters given conditions on the initial estimate (Zheng and Lafferty, , 2016; Zhang et al., , 2023). Details of FGD are given in Algorithm 1.
@@
1:Gradients of the loss function $\nabla_{U}\operatorname*{\mathcal{L}}$ and $\nabla_{V}\operatorname*{\mathcal{L}}$, initial estimate $\Phi_{0}$ and step size $\eta$.
@@
3:for $k=1,\ldots$do
4:    Update $U^{(k+1)}=U^{(k)}-\eta\nabla_{U}{\operatorname*{\mathcal{L}}}|_{U=U^{(k)},V=V^{(k)}}$.
5:    Update $V^{(k+1)}=V^{(k)}-\eta\nabla_{V}{\operatorname*{\mathcal{L}}}|_{U=U^{(k+1)},V=V^{(k)}}$.
@@
For the initialization of FGD, we use the unconstrained maximum likelihood estimator (MLE) of $\Phi^{*}$. As the likelihood function is convex, we can apply a convex optimization algorithm such as the gradient-based algorithm or L-BFGS (Liu and Nocedal, , 1989) to calculate the initialization.
With the estimate $\hat{\Phi}$ from Algorithm 1, we obtain the estimated subspaces $\hat{U}$ and $\hat{V}$ via SVD, $\hat{\Phi}=\hat{U}\hat{D}\hat{V}^{\top}$. Next, we rotate the original subspace with respect to the estimated subspace. We partition $U^{*},V^{*},\hat{U},\hat{V}$ into $(U^{*}_{1},U_{2}^{*}),(V_{1}^{*},V_{2}^{*}),(\hat{U}_{1},\hat{U}_{2}),(\hat{V}_{1},\hat{V}_{2})$, where $(\cdot)_{1}$ denotes the first $r$ columns and $(\cdot)_{2}$ denotes the remaining columns of a matrix. Define:
@@
We adopt the “subtraction method” prevalent in the low-rank matrix bandit (Kveton et al., , 2017; Jun et al., , 2019; Kang et al., , 2022). Note that $\hat{U}_{2}^{\top}U_{1}^{*}$ and $\hat{V}_{2}^{\top}V_{1}^{*}$ are expected to be small, assuming that $\hat{U},\hat{V}$ are good estimations of $U^{*},V^{*}$. Thus its product $(\hat{U}_{2}^{\top}U_{1}^{*})(\hat{V}_{2}^{\top}V_{1}^{*})^{\top}$ is expected to be negligible. To utilize this structure, we introduce the notation of “truncation-vectorization” of matrices. Consider $\Theta\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ with block-sub-matrices $\Theta_{11}\in\operatorname*{\mathbb{R}}^{r\times r}$, $\Theta_{12}\in\operatorname*{\mathbb{R}}^{r\times(d_{1}-r)}$, $\Theta_{21}\in\operatorname*{\mathbb{R}}^{(d_{2}-r)\times r}$, $\Theta_{22}\in\operatorname*{\mathbb{R}}^{(d_{2}-r)\times(d_{1}-r)}$. Define “truncation-vectorization” of $\Theta$ as
[…]
where “$rtv$” is an abbreviation of “rotation-truncation-vectorization” to represent that the original matrix $\Phi^{*}$ has been rotated with respect to $\hat{U}$ and $\hat{V}$, then truncated and vectorized.
@@ 3.2 UCB-based Stage
The second stage of the algorithm employs the UCB approach on $\theta^{*}_{rtv}$. UCB algorithm, also known as “optimism in the face of uncertainty” (Lattimore and Szepesvári, , 2020), operates based on the principle that the decision-maker should select the assortment with the highest “optimistic reward”, or the upper confidence bound of the estimated reward. Even if an assortment has a low estimated expected reward, when it has a high upper confidence bound due to large uncertainty, the algorithm can still favor that assortment. This approach inherently strikes a balance between exploiting the estimated reward and exploring assortments with higher uncertainty. As the algorithm is encouraged to further explore assortments with greater uncertainty, the confidence bounds gradually tighten, converging towards the true expected reward.
@@
After obtaining the optimistic utilities, this problem can be solved via the StaticMNL policy introduced for the static assortment problem (Rusmevichientong et al., , 2010).
During the construction of confidence bounds for the optimistic reward, the algorithm simultaneously updates the estimate $\widehat{\theta}_{t,rtv}$. Distinct from the rank-constrained MLE in the first stage, in the UCB-based stage we solve a new estimator for the updated parameter $\theta^{*}_{rtv}$. Since the parameter dimensionality has reduced from $d_{1}\times d_{2}$ in the first stage to $(d_{1}+d_{2})r-r^{2}$ in this UCB-stage, we can calculate the MLE directly on the reduced space to obtain $\widehat{\theta}_{n,rtv}=\operatorname*{arg\,min}\operatorname*{\mathcal{L}}_{n,rtv}(\theta_{rtv})$, where $\operatorname*{\mathcal{L}}_{n,rtv}$ is the negative log-likelihood in the reduced space with
@@ 4.1 Assumptions
The boundness of feature vectors is common in contextual bandit and assortment problems (Li et al., , 2017; Oh and Iyengar, , 2021), and the invertibility of the second moment ensures that the distribution of user features is sufficient to estimate the related parameters. In addition, note that we only require i.i.d. condition on the user features, not on the joint user and item features as in existing dynamic assortment literature (Oh and Iyengar, , 2021; Chen et al., , 2020).
@@
We assume that item features are also bounded with an postive definite second moment. Note that the postive definite condition is needed to guarantee the identifiability of the matrix parameter $\Phi^{*}$, and can be verified in practice since $p_{i}$’s are given in advance. Next, we impose a standard assumption in contextual MNL problems (Oh and Iyengar, , 2021), which is a slight modification of the assumption on the link function in generalized linear contextual bandits (Li et al., , 2017).

### v2

@@ 2 Problem Formulation
In this section, we introduce the mathematical formulation of the LOAD problem. At time $t$, a user with feature vector $q_{t}\in\operatorname*{\mathbb{R}}^{d_{1}}$ arrives at the platform. Assume there are $N$ products, each with item feature vector $p_{i}\in\operatorname*{\mathbb{R}}^{d_{2}}$ for $i\in[N]=\{1,\ldots,N\}$. The platform offers an assortment $S_{t}$ of size at most $K$ from the catalog $[N]$. In other words, $S_{t}\subset[N]$ and $|S_{t}|\leq K$. When the user is provided with an assortment, the user either chooses one of the items in the assortment or chooses not to purchase at all. To quantify this user choice, let $y_{it}\in\{0,1\}$ represent whether user $t$ chooses item $i$, for $i\in[N]$ and $t\geq 1$. In addition, denote $y_{0t}=1$ as the indicator that user $t$ does not choose any item at all. Let $i_{t}\in S_{t}$ be the item that user $t$ chooses (0 if no purchase). Note that for at each user $t$, $y_{i_{t},t}=1$ and the remaining is $y_{jt}=0$ for $j\in S_{t}\cup\{0\}\backslash\{i_{t}\}$, i.e., $\sum_{i\in S_{t}\cup\{0\}}y_{it}=1$. Denote $\textbf{y}_{t}=(y_{1t},\ldots,y_{Nt})^{\top}\in\{0,1\}^{N}$ as the choice vector of user $t$.
In assortment selection, a widely adopted choice model is the multinomial logit (MNL) model (McFadden, 1974; Rusmevichientong et al., 2010; Agrawal et al., 2019). Suppose the utility of item $i$ to user $t$ is given as a random variable $X_{it}=v_{it}+\epsilon_{it}$, where $v_{it}$ is the expected utility and $\epsilon_{it}$ is a mean-zero error term. Assume $v_{0t}=0$ for identifiability. The MNL model assumes that the probability of user $t$ choosing item $i$ is given as
@@
where $\Phi^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ is a rank-$r$ matrix, illustrated in Figure 4, with singular values $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$.
After the user’s purchasing decision under the MNL model, the platform earns revenue from the selected item (or zero if no purchase is made). Let $r_{i}$ denote the revenue that the platform achieves when item $i$ is sold. Given an assortment $S$, the revenue that the platform achieves from user $t$ is $\sum_{i\in S}r_{i}y_{it}$, and the expected revenue of the platform for user $t$ is
@@
The platform aims to choose the personalized assortment for each user $t$ at time $t$ to maximize the cumulative expected revenue gained from all users. Note that the possible assortment set with constraint $K$, i.e., $\operatorname*{\mathcal{S}}=\{S\subset[N]:|S|\leq K\}$ is finite. Therefore, when $\Phi^{*}$ is known in hindsight, there exists an optimal assortment $S_{t}^{*}$ for user $t$ that maximizes the expected revenue since we can calculate the expected utility for every $S\in\operatorname*{\mathcal{S}}$. Denote the optimal assortment at time $t$ as
@@
Note that maximizing the cumulative revenue is equivalent to minimizing the cumulative regret. At time $t$, given the current user context $q_{t}$ and historical data, including past assortments $S_{1},S_{2},\ldots,S_{t-1}$ presented to users $1,2,\ldots,t-1$ with respective features $q_{1},q_{2},\ldots,q_{t-1}$, and the corresponding feedback $\textbf{y}_{1},\textbf{y}_{2},\ldots,\textbf{y}_{t-1}$, the platform needs to determine the assortment $S_{t}$ for the user $t$ so that the cumulative regret could be small.
@@ 3.1 Estimation of Subspace
The first stage of the algorithm aims to acquire estimations of low-rank subspaces. Let $\Phi^{*}=U^{*}D^{*}(V^{*})^{\top}$ be the singular value decomposition of $\Phi^{*}$, where $U^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{2}},V^{*}\in\operatorname*{\mathbb{R}}^{d_{1}\times d_{1}}$ are square orthonormal matrices and $D^{*}\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ is a diagonal matrix with singular values as its diagonal components. Let $u^{*}_{(j)}$ be the $j$-th column of $U^{*}$ and $v^{*}_{(j)}$ be the $j$-th column of $V^{*}$. Consider dual contexts $p\in\operatorname*{\mathbb{R}}^{d_{2}}$ and $q\in\operatorname*{\mathbb{R}}^{d_{1}}$. The bilinear form of $p$ and $q$ with respect to $\Phi^{*}$ can be rewritten as
@@
where $\operatorname*{\mathcal{L}}_{n}(\Phi)$ is the negative log-likelihood of the MNL choice model with $n$ samples, expressed as
@@
To solve this rank-constrained optimization, we use the Burer-Monteiro formulation (Burer and Monteiro, 2003) to optimize over $U\in\operatorname*{\mathbb{R}}^{d_{2}\times r}$ and $V\in\operatorname*{\mathbb{R}}^{d_{1}\times r}$, where $\Phi=UV^{\top}$. We add a regularization term to enforce identifiability (Zheng and Lafferty, 2016). This leads to our final optimization problem:
@@
$U\leftarrow U-\eta\cdot\nabla_{U}\tilde{\operatorname*{\mathcal{L}}}_{n}(U,V),\qquad V\leftarrow V-\eta\cdot\nabla_{V}\tilde{\operatorname*{\mathcal{L}}}_{n}(U,V),$
where $\eta$ is the learning rate, $\nabla_{U},\nabla_{V}$ are gradients with respect to each component when other parameters are fixed. Applying these gradient descent steps alternatively, the estimated parameters are known to converge to the true parameters given conditions on the initial estimate (Zheng and Lafferty, 2016; Zhang et al., 2023). Details of FGD are given in Algorithm 1.
@@
1:Gradients of the loss function $\nabla_{U}\tilde{\operatorname*{\mathcal{L}}}$ and $\nabla_{V}\tilde{\operatorname*{\mathcal{L}}}$, initial estimate $\Phi_{0}$ and step size $\eta$.
@@
3:for $k=1,\ldots$, do
4:  Update $U^{(k+1)}=U^{(k)}-\eta\nabla_{U}{\tilde{\operatorname*{\mathcal{L}}}}|_{U=U^{(k)},V=V^{(k)}}$.
5:  Update $V^{(k+1)}=V^{(k)}-\eta\nabla_{V}{\tilde{\operatorname*{\mathcal{L}}}}|_{U=U^{(k+1)},V=V^{(k)}}$.
@@
For the initialization of FGD, we use the unconstrained maximum likelihood estimator (MLE) of $\Phi^{*}$. As the likelihood function is convex, we can apply a convex optimization algorithm such as the gradient-based algorithm or L-BFGS (Liu and Nocedal, 1989) to calculate the initialization.
With the estimate $\hat{\Phi}$ from Algorithm 1, we obtain the estimated subspaces $\hat{U}$ and $\hat{V}$ via SVD, $\hat{\Phi}=\hat{U}\hat{D}\hat{V}^{\top}$. Next, we rotate the original subspace with respect to the estimated subspace. We partition $U^{*},V^{*},\hat{U},\hat{V}$ into $(U^{*}_{1},U_{2}^{*}),(V_{1}^{*},V_{2}^{*}),(\hat{U}_{1},\hat{U}_{2}),(\hat{V}_{1},\hat{V}_{2})$, where $(\cdot)_{1}$ denotes the first $r$ columns and $(\cdot)_{2}$ denotes the remaining columns of a matrix. Define
@@
We adopt the “subtraction method” prevalent in the low-rank matrix bandit (Kveton et al., 2017; Jun et al., 2019; Kang et al., 2022). Note that $\hat{U}_{2}^{\top}U_{1}^{*}$ and $\hat{V}_{2}^{\top}V_{1}^{*}$ are expected to be small, assuming that $\hat{U},\hat{V}$ are good estimations of $U^{*},V^{*}$. Thus its product $(\hat{U}_{2}^{\top}U_{1}^{*})(\hat{V}_{2}^{\top}V_{1}^{*})^{\top}$ is expected to be negligible. To utilize this structure, we introduce the notation of “truncation-vectorization” of matrices. Consider $\Theta\in\operatorname*{\mathbb{R}}^{d_{2}\times d_{1}}$ with block-sub-matrices $\Theta_{11}\in\operatorname*{\mathbb{R}}^{r\times r}$, $\Theta_{12}\in\operatorname*{\mathbb{R}}^{r\times(d_{1}-r)}$, $\Theta_{21}\in\operatorname*{\mathbb{R}}^{(d_{2}-r)\times r}$, $\Theta_{22}\in\operatorname*{\mathbb{R}}^{(d_{2}-r)\times(d_{1}-r)}$. Define “truncation-vectorization” of $\Theta$ as
$\theta_{rtv}:=(\operatorname*{\textup{vec}}(\Theta_{11})^{\top},\operatorname*{\textup{vec}}(\Theta_{12})^{\top},\operatorname*{\textup{vec}}(\Theta_{21})^{\top})^{\top}\in{\operatorname*{\mathbb{R}}}^{(d_{1}+d_{2})r-r^{2}},$ (3)
where “rtvrtv” is an abbreviation of “rotation-truncation-vectorization” to represent that the original matrix $\Phi^{*}$ has been rotated with respect to $\hat{U}$ and $\hat{V}$, then truncated and vectorized.
@@
When the rank $r$ of the parameter matrix is unknown, we can estimate the rank using Generalized Information Criterion (GIC). GIC (Konishi and Kitagawa, 1996; Fan and Tang, 2013; Morimoto et al., 2024; Park et al., 2024) is a generalization of information criteria including Bayesian Information Criterion (BIC) and Akaike Information Criterion (AIC) (Akaike, 1987), and is widely used in the high-dimensional estimation problems to select hyper parameters to avoid over-fitting.
Adapting to our setting, we utilize the GIC in the following form:
$GIC(r):={\operatorname*{\mathcal{L}}}_{n}(\hat{\Phi}_{r})+a_{n}\cdot(d_{1}+d_{2}-r)\cdot r,$ (4)
which penalizes the number of free parameters under the low-rank assumption. We show that the selection of $r$ is theoretically consistent with proper selection of $a_{n}$ (Proposition 2 in the Appendix) and demonstrate the accuracy of rank selection with $a_{n}=\log(n)/n$ in our numerical results (Figure 5). We provide details of incorporating GIC to ELSA-UCB in Appendix Section S.5.2.
@@ 3.2 UCB-based Stage
The second stage of the algorithm employs the UCB approach on $\theta^{*}_{rtv}$. UCB algorithm, also known as “optimism in the face of uncertainty” (Lattimore and Szepesvári, 2020), operates based on the principle that the decision-maker should select the assortment with the highest “optimistic reward”, or the upper confidence bound of the estimated reward. Even if an assortment has a low estimated expected reward, when it has a high upper confidence bound due to large uncertainty, the algorithm can still favor that assortment. This approach inherently strikes a balance between exploiting the estimated reward and exploring assortments with higher uncertainty. As the algorithm is encouraged to further explore assortments with greater uncertainty, the confidence bounds gradually tighten, converging towards the true expected reward.
@@
After obtaining the optimistic utilities, this problem can be solved via the StaticMNL policy introduced for the static assortment problem (Rusmevichientong et al., 2010).
During the construction of confidence bounds for the optimistic reward, the algorithm simultaneously updates the estimate $\widehat{\theta}_{t,rtv}$. Distinct from the rank-constrained MLE in the first stage, in the UCB-based stage, we solve a new estimator for the updated parameter $\theta^{*}_{rtv}$. Since the parameter dimensionality has reduced from $d_{1}\times d_{2}$ in the first stage to $(d_{1}+d_{2})r-r^{2}$ in this UCB-stage, we can calculate the MLE directly on the reduced space to obtain $\widehat{\theta}_{n,rtv}=\operatorname*{arg\,min}\operatorname*{\mathcal{L}}_{n,rtv}(\theta_{rtv})$, where $\operatorname*{\mathcal{L}}_{n,rtv}$ is the negative log-likelihood in the reduced space with
@@
Remark 1 (Cold-start and Long-tail Issues).
A common challenge in recommender systems is the cold-start problem (Schein et al., 2002; Lika et al., 2014), where either new users or new items arrive with limited or no interaction history. Our framework addresses these scenarios through its contextual and low-rank structure. Each arriving user is represented by a feature vector, so the model does not depend on repeat arrivals. Similarly, when a new item is introduced, its feature vector $p_{i}$ can be directly incorporated into the bilinear utility model $p_{i}^{\top}\Theta^{*}q_{t}$, enabling preference estimation and uncertainty quantification without requiring past interaction data. This ensures robustness in environments with dynamically evolving user and item sets.
For long-tail items that may be chosen infrequently, the bilinear low-rank formulation helps mitigate data sparsity by sharing statistical strength across items and users. This allows utilities to be estimated reliably even when direct observations of a particular item are limited. Nonetheless, in cases where item or user features are partially unobserved, imputations based on historical data or clustering methods can be employed, though such procedures may introduce bias. Additional strategies, such as incorporating richer content-based features or adopting active exploration targeted at under-represented items, may further improve performance, and we view these extensions as promising directions for future research.
@@ 4.1 Assumptions
The boundness of feature vectors is common in contextual bandit and assortment problems (Li et al., 2017; Oh and Iyengar, 2021), and the invertibility of the second moment ensures that the distribution of user features is sufficient to estimate the related parameters. In addition, note that we only require the i.i.d. condition on the user features, not on the joint user and item features as in the existing dynamic assortment literature (Oh and Iyengar, 2021; Chen et al., 2020).
@@
We assume that item features are also bounded with a positive definite second moment. Note that the positive definite condition is needed to guarantee the identifiability of the matrix parameter $\Phi^{*}$, and can be verified in practice since $p_{i}$’s are given in advance. Next, we impose a standard assumption in contextual MNL problems (Oh and Iyengar, 2021), which is a slight modification of the assumption on the link function in generalized linear contextual bandits (Li et al., 2017).
@@
This assumption ensures that the choice probabilities under the MNL model are “well-behaved” around the true parameter $\Phi^{*}$. Intuitively, this assumption prevents degeneracy in the choice probabilities. That is, it avoids situations where one item dominates the entire choice set (i.e., is almost always chosen), or where the no-purchase probability becomes negligible or overwhelming. Such extreme cases would make the model highly sensitive to noise and impede reliable estimation of user preferences. In MNL-bandit models, analogous assumptions have been used in prior work, including Cheung and Simchi-Levi (2017), Chen et al. (2020), and Oh and Iyengar (2021).